# TempoPFN: Synthetic Pre-training of Linear RNNs for Zero-shot Time Series Forecasting

### v2

@@ 3.3 Data Augmentations
Augmentation Pipeline. The offline pipeline applies transformations in a structured sequence. Base series undergo optional normalization (80% probability) using random scalers (Robust, MinMax, Median, or Mean). Early-stage TS-Mixup (Darlow et al., 2023) creates convex combinations of multiple source series with probability $p=0.5$. The core augmentation step samples 2-5 distinct transformation categories with weighted probabilities: Invariances (0.6), Structure (0.6), Seasonality (0.5), Signal Processing (0.4), Discrete Effects (0.6), and Measurement Artifacts (0.3). From each selected category, one specific transformation is randomly chosen and applied in fixed global order. Optional stochastic convolution filtering (probability $p=0.3$) applies 1-3 random 1D convolutions with randomized parameters. Late-stage TS-Mixup provides additional combination opportunities, followed by finishing transformations, including minor global scaling and low-magnitude Gaussian noise injection In the following, we provide details on the augmentations we implemented.
@@
Seasonal effects simulate real-world periodicities through calendar injections that apply multiplicative factors for weekend dips, month-end spikes, and holiday-like impulses using timestamp metadata. Amplitude modulation applies localized scaling to random segments, simulating time-varying volatility. Signal processing transformations include Gaussian smoothing followed by finite-difference operators (Sobel, Laplacian, higher-order derivatives up to 4th order) and numerical integration, with outputs rescaled to preserve original value ranges. Random convolution layers with highly randomized parameters (Dempster et al., 2020) provide additional signal transformation capabilities.
Measurement artifacts introduce realistic data collection imperfections: censoring clips values at random quantiles (similarly used by TiRex (Auer et al., 2025)), non-uniform quantization maps values to discrete levels using quasi-random Sobol sequences, and resampling artifacts downsample and upsample series with various interpolation methods.
Combination strategies We implement TS-Mixup (Ansari et al., 2024) to generate novel series through convex combinations of 2-10 source series, with mixing weights sampled from Dirichlet distributions and extend it with time-dependent mixing using smooth simplex path interpolation.
Figure 4: Comparison of TempoPFN performance against other models on GiftEval benchmark. We compute both normalized and average ranks for CRPS and MASE. Colors represent the class of time series model.
@@ 4 Experiments
Pre-training Setup. TempoPFN’s pre-training is conducted exclusively on synthetic data, ensuring no exposure to real-world benchmarks prior to evaluation. The training corpus consists of approximately 10 million time series from our generators, each with a maximum length of 2048. We train our main model (34.69M parameters) using a two-stage protocol: 3 million iterations at peak LR $2\times 10^{-4}$, followed by 2 million iterations at $2\times 10^{-5}$, using the AdamW optimizer (Loshchilov & Hutter, 2019) and quantile loss. We selected a 35M model (10 layers, 4 heads, 512 embedding dimension) for its strong performance and comparability to TiRex (Auer et al., 2025). To ensure robustness across sequence lengths, we randomly sample both the context length and historical window size during training. Complete training details are in Section A.1.
@@
Figure 6: Normalized CRPS (relative to TempoPFN’s CRPS at 0% NaNs) of TempoPFN and TiRex as a function of the percentage of missing values (NaN) in the data.
Quantitative Results. We evaluate TempoPFN on the Gift-Eval benchmark, a comprehensive zero-shot forecasting suite covering diverse real-world datasets across domains and horizons. TempoPFN surpasses probabilistic performance of TabPFN-TS, the strongest synthetic-only baseline, with an overall CRPS of 0.537 (vs. 0.544). Its point-forecast accuracy is competitive, though slightly lower, with an overall MASE of 0.797 (vs. 0.771). Remarkably, despite relying solely on synthetic training data, TempoPFN matches or exceeds several leading models trained on real-data, including Chronos Bolt B (0.574/0.808), TimesFM 2.0 (0.550/0.758), and YingLong 50m (0.567/0.822), ranking 6th overall in CRPS and 5th in MASE. Figure 4 summarizes these quantitative comparisons against state-of-the-art baselines.
Qualitative Results. Figure 5 shows forecasting results on representative Gift-Eval series with varying temporal patterns (see Figure 18 for additional results). All models capture key trends and seasonality, but TempoPFN produces coherent predictive distributions without artifacts. Compared to TabPFN-TS, TempoPFN generates smoother uncertainty bounds while maintaining competitive point forecasts. This is likely a result of TabPFN-TS predicting all future time-steps in isolation while our architecture allows future time-steps to communicate. In many longer predictions made by TiRex (e.g. bizitops service), we find high frequency artifacts in the prediction of the quantiles which we hypothesize to be a result of the windowing done by TiRex which compresses the time-series into chunks of size 32 before applying the model and later upprojects them back to the original resolution. Since TempoPFN requires no windowing, we did not notice similar artifacts.
@@ 4.1 The Importance of TempoPFN’s Components
Ablation Study Setup. To manage the significant computational cost of pre-training, not all ablation experiments were conducted for a full 4M-iteration training schedule. Therefore, these studies are designed to provide strong directional evidence on the relative importance of each component, rather than to measure their full, converged performance impact.
@@
Table 2: The impact of data augmentation on model performance, shown by normalized performance values. Individual augmentation effects and probability tuning were not exhaustively explored due to resource constraints.
Gift-ZS Overall Gift-ZS Short Gift-ZS Medium Gift-ZS Long
Model CRPS MASE CRPS MASE CRPS MASE CRPS MASE
w/ Aug 0.577 0.842 0.563 0.763 0.566 0.900 0.631 1.019
w/o Aug 0.610 0.875 0.582 0.783 0.617 0.963 0.643 1.059
seasonal_naive 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Ablating the augmentation pipeline. To quantify the impact of augmentations, we trained models with and without the full suite augmentation suite (see Table 2). Results on the GIFT-Eval benchmark reveal consistent gains from our augmentation pipeline yielding a  5.4% relative improvement in overall CRPS and a  3.8% improvement in overall MASE. These gains are present across all forecasting horizons, underscoring the benefit of our complex data augmentation pipeline for extrapolation to real-world data.
Table 3: Ablation study of synthetic data priors using a leave-one-out methodology. The ’Ablation’ column indicates the single prior excluded from the training mixture. Performance is measured by CRPS and MASE (lower is better). Rows are colored to indicate the performance impact on the overall CRPS when a prior is removed: High Impact ($25\%$ increase) and Medium Impact ($10\%$ increase). N = Novel prior (our contribution), A = Adapted from open-source.
@@ 5 Conclusion and Future Work
We introduce TempoPFN, a novel time series foundation model demonstrating that linear RNNs, specifically the GatedDeltaProduct architecture, provide a highly efficient and scalable solution for zero-shot forecasting. By enabling parallelizable training, our model processes long sequences without patching or summarization heuristics. TempoPFN is trained exclusively on our open-source synthetic data generation pipeline, which integrates diverse generators and a complex augmentation framework. This synthetic-only approach ensures full reproducibility and eliminates data leakage concerns. On the Gift-Eval benchmark, TempoPFN achieves top-tier competitive performance, surpassing other synthetic-only approaches and the vast majority of models trained on real-world data, establishing linear RNNs as a powerful and scalable alternative to prevailing architectures.
@@ Acknowledgements
Appendix A Details on Synthetic Data Generation

### v3

@@ 3.3 Data Augmentations
Augmentation Pipeline. The offline pipeline applies transformations in a structured sequence. Base series undergo optional normalization (80% probability) using random scalers (Robust, MinMax, Median, or Mean). Early-stage TS-Mixup (darlow2023tsmix) creates convex combinations of multiple source series with probability $p=0.5$. The core augmentation step samples 2-5 distinct transformation categories with weighted probabilities: Invariances (0.6), Structure (0.6), Seasonality (0.5), Signal Processing (0.4), Discrete Effects (0.6), and Measurement Artifacts (0.3). From each selected category, one specific transformation is randomly chosen and applied in fixed global order. Optional stochastic convolution filtering (probability $p=0.3$) applies 1-3 random 1D convolutions with randomized parameters. Late-stage TS-Mixup provides additional combination opportunities, followed by finishing transformations, including minor global scaling and low-magnitude Gaussian noise injection. In the following, we provide details on the augmentations we implemented.
@@
Seasonal effects simulate real-world periodicities through calendar injections that apply multiplicative factors for weekend dips, month-end spikes, and holiday-like impulses using timestamp metadata. Amplitude modulation applies localized scaling to random segments, simulating time-varying volatility. Signal processing transformations include Gaussian smoothing followed by finite-difference operators (Sobel, Laplacian, higher-order derivatives up to 4th order) and numerical integration, with outputs rescaled to preserve original value ranges. Random convolution layers with highly randomized parameters (dempster2020rocket) provide additional signal transformation capabilities.
Measurement artifacts introduce realistic data collection imperfections: censoring clips values at random quantiles (similarly used by TiRex (auer2025tirex)), non-uniform quantization maps values to discrete levels using quasi-random Sobol sequences, and resampling artifacts downsample and upsample series with various interpolation methods.
Combination strategies. We implement TS-Mixup (ansari2024chronos) to generate novel series through convex combinations of 2-10 source series, with mixing weights sampled from Dirichlet distributions and extend it with time-dependent mixing using smooth simplex path interpolation.
Figure 4: Comparison of TempoPFN (3M iterations with a per-GPU batch size of 40) against other models on GiftEval benchmark. We compute both normalized and average ranks for CRPS and MASE. Colors represent the class of time series model.
@@ 4 Experiments
We ran experiments on three standard zero-shot forecasting benchmarks: Gift-Eval aksu2024gifteval, Chronos-ZS (ansari2024chronos), and fev-bench (shchur2025fev).
Pretraining Setup. TempoPFN’s pretraining is conducted exclusively on synthetic data, ensuring no exposure to real-world benchmarks prior to evaluation. The training corpus consists of approximately 10 million time series from our generators, each with a maximum length of 2048. We train our main model with the AdamW optimizer (loshchilov-iclr19a) to minimize the quantile loss, for a total of 3 million iterations and a batch size of 40. We selected a 40M model (10 layers, 4 heads, 512 embedding dimension) for its strong performance and comparability to TiRex (auer2025tirex). To ensure robustness across sequence lengths, we randomly sample both the context length and historical window size during training. Complete training details are in Appendix D.
4.1 Results on Gift-Eval
Quantitative Results. We evaluate TempoPFN on the Gift-Eval benchmark (aksu2024gifteval), a comprehensive zero-shot forecasting suite covering 23 diverse real-world datasets across domains and horizons. TempoPFN surpasses probabilistic performance of TabPFN-TS, the strongest synthetic-only baseline, with an overall CRPS of 0.537 (vs. 0.544). Its point-forecast accuracy is competitive, though slightly lower, with an overall MASE of 0.797 (vs. 0.771). Remarkably, despite relying solely on synthetic training data, TempoPFN matches or exceeds several leading models trained on real-data, including Chronos Bolt B (0.574/0.808), TimesFM 2.0 (0.550/0.758), and YingLong 50m (0.567/0.822), ranking 6th overall in CRPS and 5th in MASE. Figure 4 summarizes these quantitative comparisons against state-of-the-art baselines.
@@ 4.1 Results on Gift-Eval
Qualitative Results. Figure 5 shows forecasting results on representative Gift-Eval series with varying temporal patterns (see Figure 22 for additional results). All models capture key trends and seasonality, but TempoPFN produces coherent predictive distributions without artifacts. Compared to TabPFN-TS, TempoPFN generates smoother uncertainty bounds while maintaining competitive point forecasts. This is likely because TabPFN-TS predicts all future time steps in isolation, whereas our architecture allows future time steps to communicate. In many longer predictions made by TiRex (e.g. bizitops service), we find high-frequency artifacts in the prediction of the quantiles, which we hypothesize to be a result of the windowing done by TiRex which compresses the time series into chunks of size 32 before applying the model and later up-projects them back to the original resolution. Since TempoPFN requires no windowing, we did not notice similar artifacts.
Figure 6: Normalized CRPS of TempoPFN and TiRex as a function of missing values (NaN) in the data.
@@
4.2 Results on Chronos-ZS
We evaluate TempoPFN on the Chronos-ZS benchmark (ansari2024chronos), comprising 27 diverse datasets from the GluonTS (gluonts) and Monash (godahewa-neuripsdbt21a) repositories, spanning multiple domains (e.g., energy, transport, healthcare) and frequencies. Figure 7 shows the aggregated performance in terms of average rank for both probabilistic (CRPS) and point (MASE) forecasting, where TempoPFN ranks again among the top-performing zero-shot models. In Figure 19 in the appendix, we also show the normalized CRPS and MASE scores on this benchmark.
Figure 7: CRPS and MASE average rank of TempoPFN (3M iterations with a per-GPU batch size of 40) on Chronos-ZS benchmark.
4.3 Results on fev-bench
We also evaluate on the recent fev-bench (shchur2025fev), which standardizes evaluation across 100 diverse forecasting tasks, as well as rigorously tracks data leakage and failure rates, making this benchmark more challenging than Gift-Eval and Chronos-ZS. fev-bench evaluates the model performance using two aggregate scores derived from the pairwise error matrices: (1) Win Rate (%), representing the fraction of model pairs and tasks where the model achieves a lower error than the competitor; and (2) Skill Score (%), a robust measure of relative error reduction compared to the Seasonal Naive baseline.
Leaderboard Results. Table 2 presents the leaderboard results for MASE, where TempoPFN is ranked sixth. We can see a similar rank in Table 13 of Appendix E.2 for the Scaled Quantile Loss (SQL), which captures calibration quality by evaluating the quality of the entire predictive distribution at each time step. Most importantly, in both metrics, our model outperforms the other leading synthetic-only baseline, TabPFN-TS (Rank 8 in both). To also visualize relative strengths in probabilistic forecasting, we show in Figure 20 in the appendix the head-to-head Win Rates and Skill Scores.
Table 2: fev-bench leaderboard based on MASE. Models are ranked by Win Rate and Skill Score.
Rank Model Avg. Win Rate (%) Skill Score (%) Median Runtime (s) Leakage (%) Failed Tasks (%) Organization Zero-shot
1 Chronos-2 88.0 35.5 3.57 0 0 AWS
2 TiRex 76.7 30.0 1.4 1 0 NX-AI
3 TimesFM 2.5 74.9 30.2 10.89 10 0 Google
4 Toto 1.0 66.5 28.2 77.51 8 0 Datadog
5 Moirai 2.0 60.5 27.3 1.9 28 0 Salesforce
6 TempoPFN 60.5 25.1 8.57 0 0 Uni Freiburg
7 Chronos-Bolt 60.1 26.5 1.0 0 0 AWS
8 TabPFN-TS 58.2 27.6 300.57 0 2 Prior Labs
9 Sundial-Base 52.4 24.7 33.99 1 0 Tsinghua University
10 Stat. Ensemble 47.1 15.7 624.45 0 11
11 AutoARIMA 35.6 11.2 120.16 0 10
12 AutoTheta 33.6 11.0 9.27 0 0
13 AutoETS 32.6 2.3 16.24 0 3
14 Seasonal Naive 20.0 0.0 2.32 0 0
@@ 4.4 The Importance of TempoPFN’s Components
We conduct several ablation studies on TempoPFN’s building blocks. Due to significant pretraining costs, not all ablation experiments were conducted on the full 3M-iteration training schedule.
@@
Table 3: Ablation study of synthetic data priors using a leave-one-out methodology (500k iterations). The ’Ablation’ column indicates the single prior excluded from the training mixture. Performance is measured by CRPS and MASE (lower is better). Rows are colored to indicate the performance impact on the overall CRPS when a prior is removed: High Impact ($25\%$ increase) and Medium Impact ($10\%$ increase). N = Novel prior (our contribution), A = Adapted from open-source.
@@
Table 4: The impact of data augmentation on model performance, shown by normalized performance values (500k iterations). Individual augmentation effects and probability tuning were not exhaustively explored due to resource constraints.
Gift-ZS Overall Gift-ZS Short Gift-ZS Medium Gift-ZS Long
Model CRPS MASE CRPS MASE CRPS MASE CRPS MASE
w/ Aug 0.577 0.842 0.563 0.763 0.566 0.900 0.631 1.019
w/o Aug 0.610 0.875 0.582 0.783 0.617 0.963 0.643 1.059
seasonal_naive 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Ablating the augmentation pipeline. To quantify the impact of augmentations, we trained models with and without the full augmentation suite (see Table 4). Results on the GIFT-Eval benchmark reveal consistent gains from our augmentation pipeline, yielding a  5.4% relative improvement in overall CRPS and a  3.8% improvement in overall MASE. These gains are present across all forecasting horizons, underscoring the benefit of our complex data augmentation pipeline for extrapolation to real-world data.
@@ 5 Conclusion and Future Work
We introduce TempoPFN, a novel time series foundation model demonstrating that linear RNNs, specifically the GatedDeltaProduct architecture, provide a highly efficient and scalable solution for zero-shot forecasting. By enabling parallelizable training, our model processes long sequences without patching or summarization heuristics. TempoPFN is trained exclusively on our open-source synthetic data generation pipeline, which integrates diverse generators and a complex augmentation framework. This synthetic-only approach ensures full reproducibility and eliminates data leakage concerns. On the Gift-Eval, fev-bench and Chronos-ZS benchmarks, TempoPFN achieves top-tier competitive performance, surpassing other synthetic-only approaches and the vast majority of models trained on real-world data, establishing linear RNNs as a powerful and scalable alternative to prevailing architectures.
@@ Acknowledgements
Appendix A Details on the Design Principles of TempoPFN
A.1 Background: Prior Data Fitted Networks (PFNs)
Prior Data Fitted Networks (PFNs) (muller2022transformers) represent a paradigm shift in machine learning, moving from learning a single fixed task to learning a universal inference algorithm. In this section, we provide a brief overview of the PFN framework to contextualize the methodology used in TempoPFN.
Definition and Objective.
A PFN is a neural network $\phi$, with parameters $\theta$, trained to approximate the posterior predictive distribution (PPD) induced by a prior distribution $P(\mathcal{D})$ over datasets. Formally, let a dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ be drawn from a prior $P$. The goal of a PFN is to minimize the Kullback-Leibler divergence (or equivalently, the cross-entropy loss) between its output and the true posterior predictive distribution of the prior:
$\min_{\theta}\mathbb{E}_{\mathcal{D}\sim P}\left[\sum_{i=1}^{N}-\log p_{\theta}(y_{i}\mid x_{i},\mathcal{D}_{1:i-1})\right]$ (2)
where $\mathcal{D}_{1:i-1}$ represents the context (history) observed so far.
In-Context Learning as Bayesian Inference.
Unlike traditional supervised learning, where the model’s weights $\theta$ encode the solution to a specific task (e.g., ”predict sales for Company X”), a PFN’s weights encode the algorithm for solving a class of tasks defined by the prior. At inference time, the PFN performs In-Context Learning (ICL): it takes a small dataset of context observations (the history of a time series) and produces predictions for new query points (the future) in a single forward pass. Crucially, this forward pass acts as a fast approximation of Bayesian Inference without the computational cost of MCMC or variational methods (muller2022transformers; hollmann2023tabpfn).
The Role of Synthetic Data.
The performance of a PFN is fundamentally bounded by the quality and diversity of its prior $P$. Since real-world data is often limited, biased, or private, PFNs are typically trained on vast repositories of synthetic data generated from procedural priors. For example, TabPFN (hollmann2023tabpfn) uses Structural Causal Models (SCMs) to generate tabular data, while ForecastPFN (dooley2023forecastpfn) uses a mix of trends and seasonalities. Similarly, in TempoPFN the ”Prior” is the novel synthetic data pipeline detailed in Appendix C, which generates diverse temporal dynamics using SDEs, GPs, and asymmetric waveforms. Our ”Network” is a Linear RNN, namely GatedDeltaProduct (siems2025deltaproduct), chosen for its efficiency in handling long sequential contexts compared to the Transformers used in previous PFNs. By training the sequence model on this synthetic prior, TempoPFN learns to infer the underlying generative process of any unseen time series given its history, enabling zero-shot probabilistic forecasting.
A.2 Design Principles for Synthetic Data Generation
Our selection of synthetic data generators is grounded in the principle of structural decomposition. We posit that the manifold of real-world time series can be spanned by four fundamental dynamical properties: Smoothness, Stochastic Volatility (Roughness), Temporal Asymmetry, and Discontinuities. Existing synthetic pipelines often over-index on the first (trends/seasonality) (dooley2023forecastpfn) while neglecting the latter three. We designed a principled portfolio of generators to act as orthogonal ”basis functions” for these properties, ensuring our prior distribution covers the complex dynamics found in downstream tasks.
Smooth Dynamics (Gaussian Processes). To capture non-parametric trends and local correlations, we employ Gaussian Processes (GPs). Real-world data is dominated by latent trends that evolve smoothly but unpredictably, such as demographic shifts or climate warming, which cannot be captured by rigid linear or polynomial functions. GPs with RBF or Matérn kernels serve as the standard for modeling such smooth, differentiable manifolds, providing the model with a robust prior for interpolation and extrapolation of continuous trends.
Stochastic Volatility and Roughness (SDEs). A critical deficiency in standard synthetic pipelines is the reliance on additive Gaussian noise ($\epsilon\sim\mathcal{N}(0,1)$), which implies homoscedasticity (constant variance). However, financial and physical systems are inherently heteroscedastic, exhibiting state-dependent volatility and mean-reverting dynamics. To fill this gap, we integrate Stochastic Differential Equations (SDEs), specifically Regime-Switching Ornstein-Uhlenbeck (OU) processes. By explicitly modeling the diffusion term $\sigma(t,y_{t})$, we force the model to learn to distinguish between deterministic signal drift and stochastic volatility clustering, a capability essential for accurate uncertainty quantification.
Asymmetric Periodicity (Sawtooth Waveforms). Standard sinusoidal generators rely on an assumption of time-reversal symmetry (equal rise and fall times). Yet, many physical and economic processes are inherently asymmetric and irreversible: inventory levels deplete gradually and restock instantaneously; capacitors discharge rapidly. We selected the Sawtooth wave as the fundamental primitive for asymmetry. Unlike Triangle waves (symmetric) or Square waves (step functions), the Sawtooth explicitly models the gradual-rise/sharp-drop dynamic.
Discontinuities and Structural Breaks (Spikes/Steps). Finally, real-world data is rife with instantaneous regime changes: policy shifts, sensor failures, or sudden shocks. All of these violate the smoothness assumptions of GPs and SDEs. To model these structural breaks, we include explicit Step and Spike generators. Including these non-differentiable primitives ensures the model remains robust to covariate shifts and prevents the ”smearing” of distinct regimes into a single average.
By composing these four distinct dynamical behaviors, TempoPFN has access to a complex prior, allowing it to generalize in a zero-shot manner to unseen time series by identifying the governing combination of dynamics (related to the state-tracking capabilities of GatedDeltaProduct too), rather than memorizing dataset-specific statistics.
A.3 Comparison with Existing Synthetic Strategies
Freq-Synth (nochumsohn2024beyond) and TabPFN-TS (hoo2024the) are two recent methods that employ only synthetic data training as well. However, their methodologies represent fundamentally different paradigms in synthetic data generation and usage compared to TempoPFN. Freq-Synth adopts a task-specific generation strategy, hence requiring prior knowledge of the target dataset’s sampling rate to generate a custom training corpus of harmonic signals (sums of sinusoids) tailored to mitigate data scarcity for that specific task. In contrast, in TempoPFN, we pretrain a single model on a fixed, comprehensive corpus designed to marginalize over diverse temporal dynamics without requiring task-specific data generation. This distinction places TempoPFN in a unique position within the broader landscape of synthetic time series data (liu2025empowering). While most existing synthetic pre-training methods (e.g., Chronos, TimesFM) rely on standard statistical components like GPs or ARMA processes, TempoPFN explicitly expands this design space by introducing novel generators for stochastic volatility (via SDEs) and temporal asymmetry (via sawtooth waves). Interestingly, TabPFN-TS relies on cross-domain adaptation, leveraging a model pre-trained on synthetic tabular data (via structural causal models) to effectively ”feature-engineer” time series problems into tabular regression tasks, rather than learning temporal dynamics directly.
Parallel to this work, graf2025flowstate introduced FlowState, a time series foundation model that also leverages State Space Models (SSMs) for subquadratic computational efficiency. FlowState features an SSM-based encoder combined with a functional basis decoder to enable sampling-rate invariance and continuous-time modeling. While both FlowState and TempoPFN move away from Transformer-based architectures in favor of linear recurrences, our approaches differ fundamentally in their pre-training data paradigms. FlowState is pre-trained on a combination of real-world datasets (subsets of GIFT-Eval and Chronos corpora), whereas TempoPFN establishes the viability of a purely synthetic pre-training pipeline. Furthermore, while FlowState focuses on resolution adaptation, TempoPFN focuses on maximizing zero-shot generalization through a diverse synthetic prior designed and utilizing the GatedDeltaProduct for state tracking.
Appendix B Details on Architectural Mechanisms
B.1 GatedDeltaProduct Architecture and State Tracking
To overcome the expressivity limitations of diagonal linear RNNs while retaining linear-time parallel scan computation, TempoPFN uses the GatedDeltaProduct recurrence (siems2025deltaproduct). Unlike diagonal SSMs such as Mamba (gu2023mamba) or RWKV (peng2023rwkv), whose state-transition matrices are restricted to diagonal structure, GatedDeltaProduct employs a structured non-diagonal transition matrix constructed as a product of generalized Householder matrix updates. This yields a more expressive class of linear operators while keeping both training and inference efficient.
Recurrence Mechanism.
Each layer maintains a matrix-valued hidden state ${\bm{H}}_{t}\in\mathbb{R}^{d\times n}$, updated via a linear recurrence:
${\bm{H}}_{t}={\bm{A}}_{t}{\bm{H}}_{t-1}+{\bm{B}}_{t},\qquad{\bm{y}}_{t}={\bm{H}}_{t}{\bm{x}}_{t}.$ (3)
Here, $x_{t}\in\mathbb{R}^{D}$ represents the input vector at the current time step $t$, where $D$ is the input dimension (e.g., number of features). ${\bm{H}}_{t}\in\mathbb{R}^{N}$ is the hidden state vector, representing the compressed memory, where $N$ is the state dimension. ${\bm{A}}_{t}\in\mathbb{R}^{N\times N}$ is the state-transition matrix, which defines how the information from the previous hidden state evolves and is stored in the new state. DeltaProduct utilizes a dense, parameterized matrix ${\bm{A}}_{t}$, which is explicitly constrained to be near orthogonal through its initialization and parameterization scheme (a product of $n_{h}$ rank-1 Householder-like updates):
${\bm{A}}_{t}=g_{t}\prod_{j=1}^{n_{h}}\left({\bm{I}}-\beta_{t,j}{\bm{k}}_{t,j}{\bm{k}}_{t,j}^{\top}\right),$ (4)
where $g_{t}\in[0,1]$ is a forget gate, ${\bm{k}}_{t,j}$ are normalized key vectors, and $\beta_{t,j}$ are step sizes predicted from the input. The orthogonality constraint ensures that the transformation applied to the hidden state, ${\bm{H}}_{t-1}$, preserves its magnitude. Consequently, the hidden state ${\bm{H}}_{t}$ can effectively maintain its stability and information content over extended time steps, enabling robust state tracking of long-term trends and cyclical patterns in time series forecasting. Because each factor ${\bm{I}}-\beta{\bm{k}}{\bm{k}}^{\top}$ is a rank-1 update, the full matrix-vector product remains $\mathcal{O}(dn_{h})$ and is fully parallelizable via a parallel prefix scan (yang2024parallelizing). Finally, ${\bm{B}}_{t}\in\mathbb{R}^{N\times D}$ is the input weight matrix. This matrix transforms the current input $x_{t}$ and integrates it into the hidden state ${\bm{H}}_{t}$.
Gating.
The gating in GatedDeltaProduct was introduced to embed essential non-linearity in the basic linear recurrence above, therefore enhancing the model’s overall expressiveness and selective memory. After the main linear recurrence is performed, its output is processed along two parallel streams: 1) Main Stream: The result of the linear recurrence, intended for the final output; 2) Gate Stream: The same recurrent output is passed through a non-linear activation function (e.g., the SiLU or Swish function). These two streams are combined via element-wise multiplication (the gate). This operation selectively controls which parts of the recurrent output are emphasized or suppressed, mirroring the functionality of sophisticated recurrent units like the Gated Recurrent Unit (GRU) or the selectivity found in Mamba’s architecture.
State Weaving.
This mechanism was specifically designed for the overall multi-layer structure of the TempoPFN framework where multiple GatedDeltaProduct layers are stacked. Instead of simply discarding the final hidden state, the mechanism weaves temporal information across the depth of the model. Specifically, the final hidden state (${\bm{H}}_{T}$) outputted by the first GatedDeltaProduct layer in the stack is passed forward to serve as the initial hidden state (${\bm{H}}_{0}$) for the subsequent GatedDeltaProduct layer. This ensures that deeper layers do not start their recurrence from a blank slate but instead build upon the aggregated temporal state representations learned by the shallower layers. This process creates a dense flow of information across both the time dimension (via the recurrence) and the model depth (via the weaving).
State Tracking and Its Relevance in Time-Series Forecasting.
A central advantage of DeltaProduct is its ability to perform state tracking, i.e., maintaining and updating information over long sequences. Diagonal linear RNNs and SSMs (e.g., Mamba, RWKV, GLA) update each hidden dimension independently, which is efficient but severely limits expressivity: they cannot mix coordinates, cannot implement basic tracking functions such as parity or counting (grazzi-iclr25a), and their states inevitably drift toward zero due to exponential decay. DeltaProduct avoids this failure mode through negative eigenvalues present in the structured non-diagonal transitions, where each factor $({\bm{I}}-\beta{\bm{k}}{\bm{k}}^{\top})$ acts as a reflection or low-rank rotation that preserves geometry and prevents collapse. This capability is crucial for time-series forecasting, where tracking corresponds to maintaining trend and level information across long contexts. As a result, GatedDeltaProduct layers maintain trend information without attenuation, enabling stable and coherent extrapolation over extended horizons.
Appendix C Synthetic Data Implementation Details
C.1 Generator Specifications